Circular No. 6074

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

JUPITER AND COMET 1993e I. de Pater, C. Heiles, and M. Wong, University of California, Berke-

ley; and R. Maddalena, National Radio Astronomy Observatory (NRAO), Green Bank, report on 20-cm radio observations of Jupiter with the 42.7-m telescope at Green Bank and the Very Large Array (VLA) in New Mexico (both facilities of the NRAO) obtained before, during, and after the comet-1993e impacts: “At Green Bank we measured the total and linearlypolarized flux density, while at the VLA we imaged the spatial brightness distribution of the planet in all four Stokes parameters. As reported before by Kesteven et al. (IAUC 6040), Klein and Gulkis (IAUC 6042), Galopeau et al. (IAUC 6045), Smits et al. (IAUC 6047) and, most recently, by Bird et al. (IAUC 6069), the total flux density of the emission increased by ∼ 20 percent during the week of cometary impacts, and slowly decreased afterwards. The linear polarization varied by ∼ 2 percent. In addition to changes in total and linearly-polarized flux density, the 42.7-m data show that the beaming curve (variation in intensity during one Jovian period) changed during the week of cometary impacts. VLA images show that Jupiter’s brightness distribution changed drastically during the week of cometary impacts. This must be related to the observed variations in the beaming curve and changes in the ‘disk-averaged’ polarization characteristics.”

PERIODIC COMET BORRELLY (1994l) Total visual magnitude estimates: Aug. 30.45 UT, 10.8 (P. Roques,

Williams, AZ, 0.25-m reflector + CCD + yellow filter); Sept. 3.15, 11.1 (J. D. Shanklin, Cambridge, UK, 0.33-m reflector); 3.44, 10.9 (C. S. Morris, Lockwood Valley, CA, 0.26-m reflector); 4.09, 9.8 (W. Hasubick, Buchloe, Germany, 0.20-m reflector); 5.11, 9.8 (Hasubick, 25×100 binoculars); 7.12, 11.8 (G. Bakos, Ráktanya, Hungary, 0.44-m reflector).

PERIODIC COMET MACHHOLZ 2 (1994o) Total visual magnitude estimates of the second component (cf. IAUC

6066): Sept. 3.49 UT, 11.5 (Morris); 4.06, 10.5 (Hasubick, 0.20-m reflector); 4.50, 11.5 (Morris, Pine Mountain Club, CA, 0.26-m reflector); 7.11, 12.6 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector).

1994 September 8 Daniel W. E. Green